OpenStudy (anonymous):

please help me to integrate ln(x^2 + 4)

OpenStudy (anonymous):

log you integrate by parts always the same trick put \(u=\ln(x^2+4), dv=1\)

OpenStudy (anonymous):

ya i got that part, but itz later on dat i get stuck

OpenStudy (anonymous):

i gotta run maybe myininaya will help

OpenStudy (anonymous):

\[\int\limits_{}^{}\ln(x^2+4)=x \ln(x^2+4)-\int\limits_{}^{}\frac{2x^2}{x^2+4}\]

OpenStudy (anonymous):

oh ok, sure thnx

OpenStudy (anonymous):

@NotSObright : i got stuck exactly there

myininaya (myininaya):

Y\[\frac{2x^2}{x^2+4}=\frac{x^2+4+x^2+4-8}{x^2+4}=\frac{x^2+4}{x^2+4}+\frac{x^2+4}{x^2+4}-\frac{8}{x^2+4}\]

myininaya (myininaya):

oops ignore that Y that was accidental type in thingy

OpenStudy (anonymous):

thank youuuuu so much everyone!!!!!!!!

myininaya (myininaya):

Wow you got it so quickly? Really you understand?

OpenStudy (anonymous):

yep....i jst u knw sorta needed some help getting outta that 2x^2/(x^2+4)

myininaya (myininaya):

cool you could have used long division but i like the way i did it :)

myininaya (myininaya):

When the degrees of the polys on top and bottom are the same do long division (or find shortcut like I did) If the degree on top is larger do long division

OpenStudy (anonymous):

thank you!!!!!!!!!!!!

myininaya (myininaya):

Np. Peace @satellite73 is awesome too! :)

OpenStudy (anonymous):

uh...ya bt i cn make both ur answers best answers, so u knw....lol

myininaya (myininaya):

I know. :( Its all cool though.

myininaya (myininaya):

:)